# A bi-fidelity method for the uncertain Vlasov-Poisson system near quasineutrality in an asymptotic-preserving PIC framework

### v1

@@
Abstract
In this paper, we study the Vlasov-Poisson system with massless electrons (VPME) near quasineutrality and with uncertainties. Based on the idea of reformulation on the Poisson equation by [P. Degond et.al., Journal of Computational Physics, 229 (16), 2010, pp. 5630–5652], we first consider the deterministic problem and develop an efficient asymptotic-preserving particle-in-cell (AP-PIC) method to capture the quasineutral limit numerically, without resolving the discretizations subject to the small Debye length in plasma. The main challenge and difference compared to previous related works is that we consider the nonlinear Poisson in the VPME system which contains $e^{\phi}$ (with $\phi$ being the electric potential) and provide an explicit scheme. In the second part, we extend to study the uncertainty quantification (UQ) problem and develop an efficient bi-fidelity method for solving the VPME system with multidimensional random parameters, by choosing the Euler-Poisson equation as the low-fidelity model. Several numerical experiments are shown to demonstrate the asymptotic-preserving property of our deterministic solver and the effectiveness of our bi-fidelity method for solving the model with random uncertainties.
Key words. Vlasov-Poission near quasineutrality, asymptotic-preserving, particle-in-cell method, uncertainty quantification, bi-fidelity method, stochastic collocation
MSC codes. 65Y20, 35J05, 82C40, 82D10
1 Introduction
@@
[…]
Here $\varepsilon$ is the dimensionless parameter characterized by the Debye length and $f(t,\bm{x},\bm{v})$ represents the distribution of ions. This VPME system has been used in plasma physics to model ion plasma, study the formation of ion-acoustic shocks [43] and the expansion of plasma into vacuum [39], among other applications. We refer to [27] for more detailed introduction.
In real applications of plasma physics, these equations often exhibit different scales, ranging from the microscopic dynamics of individual particles to the macroscopic behavior of the plasma as a whole. When the scaled Debye length $\varepsilon\ll 1$, the plasma is called quasineutral: since the scale of charge separation is very small, the plasma appears to be neutral at the scale of observation. Quasineutrality is a common and important phenomenon for plasmas. In [11] the author emphasizes quasineutrality as one of the key properties that distinguish plasmas from ionised gases, we also refer to [25] and other theoretical works [7, 8, 9, 20, 21, 22, 23, 24, 28, 29] which give a review of recent developments on for the Vlasov-type equations near quasineutrality.
@@
Numerical computation for the VPME system (1.1), that is under study in this work, seems underdeveloped. For the linear Poisson equation, a key idea was first proposed by P. Degond et al. [12], which is to reformulate the system to a new one that remains uniformly elliptic thus can capture the quasineutral limit of the system. Later in [14], the authors developed an asymptotic-preserving PIC method for the Vlasov-Poisson system that can capture the quasineutral limit at the discrete level. Previous works on AP methods for the quasineutral limit have been devoted to the Euler-Poisson model [13, 15] and the Vlasov-Poisson problem [3]. In this work, we will develop an efficient AP-PIC method for solving the VPME system (1.1). For the deterministic problem, the main difficulty compared to the previous work [14] is that we consider the nonlinear Poisson equation with $e^{\phi}$ on the right-hand-side, thus the derivation of the quasineutral limit and the design of our discretized scheme is novel and new. Moreover, we provide a stable, explicit scheme that allows us to solve the reformulated Poisson equation efficiently.
@@
In industrial applications of plasma, it is important to study the uncertainty quantification (UQ) problems since there are many sources of uncertainties in the model. In the second part of this paper, we will extend to study UQ problems for the VPME system and consider the non-intrusive stochastic collocation (SC) method [1, 5, 41, 48] in a bi-fidelity framework. One key challenge of SC approaches is the computational cost, as they require repetitive implementations of the deterministic solver. Fortunately, in practice we can usually find some approximated low-fidelity models that are computationally cheaper but can still characterize behaviour of the complex model in the random space. The multi-fidelity algorithm [40, 49] is to combine the computational efficiency of low-fidelity models with the high accuracy of high-fidelity models, to construct an accurate surrogate for the high-fidelity model at a significantly reduced simulation cost. Regarding UQ for kinetic equations with random parameters, one can refer to a series of works using the bi-fidelity methods [4, 36, 37, 42].
@@
The rest of the paper is organized as follows. In Section 2, we first give an introduction of the Vlasov-Poisson system with massless electron and its related macroscopic equations. In Section 3, we begin with reviewing the classical PIC method, then derive the reformulated Poisson equation that leads to the design of our fully discretized AP-PIC scheme, with a summary of the method shown in Figure 1. In Section 4, we introduce the bi-fidelity method and how we choose the high-fidelity and low-fidelity model (Euler-Poisson) in solving the VPME system with uncertainties. Several numerical examples, including the deterministic and problem with random parameters have been shown in Section 5, which demonstrate the effectiveness and AP property of our solver for the VPME system, in addition to accuracy and efficiency of our designed bi-fidelity method for the corresponding UQ problems. Finally, we conclude the paper and mention some future work.
2 Models
@@
(VPME): $\partial_{t}f+\bm{v}\cdot\nabla_{\bm{x}}f-\nabla_{\bm{x}}\phi\cdot\nabla_{\bm{v}}f=0,$ (2.1a)
(VPME): $-\varepsilon^{2}\Delta_{\bm{x}}\phi=\int f\mathrm{d}\bm{v}-e^{\phi},$ (2.1b)
(VPME): $f(0,\bm{x},\bm{v})=f_{0},\,\int f_{0}\,\mathrm{d}\bm{v}\mathrm{d}\bm{x}=1.$ (2.1c)
where $f(t,\bm{x},\bm{v})$ is interpreted as the distribution of ions, at position $\bm{x}$ moving with velocity $\bm{v}$. Here $\phi$ is the electric potential that is governed by the Poisson equation, and $\varepsilon$ is the dimensionless parameter characterized by the Debye length. More details of derivation of the model can be found in [28].
@@
$f(t,\bm{x},\bm{v})=n(t,\bm{x})\delta(\bm{v}-\bm{u}(t,\bm{x})),$ (2.2)
with $\delta$ being the Dirac delta function. Here, $n(t,\bm{x})$ is the ion density and $u(t,\bm{x})$ is the ion velocity given by
$n(t,\bm{x})=\int f(t,\bm{x},\bm{v})\mathrm{d}\bm{v},\quad\bm{u}(t,\bm{x})=\frac{1}{n(t,\bm{x})}\int\bm{v}f(t,\bm{x},\bm{v})\mathrm{d}\bm{v}.$ (2.3)
Then one derives isothermal Euler system (IsE) by taking the first and second moments in velocity:
[…]
If the dimensionless electron potential satisfies $\phi\ll 1$, one can get a simplified model with linearized Maxwell-Boltzmann law,
[…]
As $\varepsilon\to 0$, by assuming the solution satisfies the form (2.2), one gets the following Euler-type equations,
$\left\{\begin{array}[]{l}\partial_{t}n+\nabla_{\bm{x}}\cdot(n\bm{u})=0,\\[6.0pt] \partial_{t}(n\bm{u})+\nabla_{\bm{x}}\cdot(n\bm{u}\otimes\bm{u})=-\frac{1}{2}\nabla_{\bm{x}}n^{2}.\end{array}\right.$ (2.6)
which can be interpreted as the shadow water system.
For the two model systems (2.1c) and (2.5), we consider periodic boundary conditions for the distribution $f(t,\bm{x},\bm{v})$ and potential $\phi$. In the one-dimensional case, the boundary conditions are given by
@@
3 AP-PIC method for VPME system
3.1 Classical PIC method
@@
3.2 Derivation of the reformulated Poisson equation
@@
$-\nabla\cdot\left((\varepsilon^{2}\partial_{tt}+n_{i})\nabla\phi\right)+\partial_{tt}n_{e}=\nabla^{2}:S_{i},$ (3.11)
@@
$\begin{array}[]{l}-\varepsilon^{2}\Delta\phi_{0}=\left(n_{i}-n_{e}\right)_{0}\\[4.0pt] -\varepsilon^{2}\Delta\phi_{0}^{\prime}=\left(n_{i}-n_{e}\right)_{0}^{\prime},\end{array}$ (3.12)
@@
where $\phi$ need to satisfy the initial data (3.12).
As $\varepsilon\to 0$, the quasineutral Vlasov system becomes
@@
We note that although the original Poisson equation (3.13) and the reformulated Poisson equation (3.14) are equivalent, they are not equally suited in the quasi-neutral limit. In fact, the quasi-neutral potential equation (3.18) automatically serves as the formal limit $\varepsilon\to 0$ of the reformulated Poisson equation (3.14), but not of the original Poisson equation (3.13). Therefore, the design of Asymptotic-Preserving schemes must be based on the reformulated system (3.17)–(3.18). We will introduce our discretized scheme in the next section.
3.3 The AP-PIC method
@@
$-\nabla\cdot\Big{(}(\Delta t)^{2}n_{i}^{m}\nabla\phi^{m+1}+\varepsilon^{2}\nabla\phi^{m+1}\Big{)}+e^{\phi^{m+1}}\\ =(\Delta t)^{2}\nabla^{2}:S_{i}^{m}+2n_{i}^{m}-n_{i}^{m-1}.$ (3.20)
@@
$-\nabla\cdot\Big{(}(\Delta t)^{2}n_{i}^{m}\nabla\phi^{m+1}\Big{)}+e^{\phi^{m+1}}=(\Delta t)^{2}\nabla^{2}:S_{i}^{m}+2n_{i}^{m}-n_{i}^{m-1}.$ (3.21)
@@
Consider the right-hand side of (3.20),the term $n_{i}^{m}-n_{i}^{m-1}$ can be replaced by $-\nabla(nu)_{i}^{m}$, we derive an equivalent form shown as
$-\nabla\cdot\Big{(}(\Delta t)^{2}n_{i}^{m}\nabla\phi^{m+1}+\varepsilon^{2}\nabla\phi^{m+1}\Big{)}+e^{\phi^{m+1}}\\ =(\Delta t)^{2}\nabla^{2}:S_{i}^{m}+n_{i}^{m}-\nabla(nu)_{i}^{m}.$ (3.22)
@@
$\begin{array}[]{l}\left(D^{+}g\right)_{j}=\frac{g_{j+1}-g_{j}}{h}\\[4.0pt] \left(D^{-}g\right)_{j}=\frac{g_{j}-g_{j-1}}{h},\\[4.0pt] \left(\Delta g\right)_{j}=\left(D^{+}D^{-}g\right)_{j}=\left(D^{-}D^{+}g\right)_{j}=\frac{g_{j+1}-2g_{j}+g_{j-1}}{h^{2}}.\end{array}$
@@
$F(\bm{\Phi}^{m+1})=A^{m}\bm{\Phi}^{m+1}+\mathrm{diag}(e^{\bm{\Phi}^{m+1}})-\bm{b}^{m}=0,$ (3.24a)
$\bm{\Phi}^{m+1}=\left[\phi^{m+1}_{1},\phi^{m+1}_{2},\dots,\phi^{m+1}_{N_{h}}\right]^{\top},$ (3.24b)
@@
The Newton iteration needs the Jacobian of $F(\bm{\Phi}^{m+1})$ defined in (3.24a), which is
$J(\bm{\Phi}^{m+1})=A^{m}+\mathrm{diag}(e^{\bm{\Phi}^{m+1}}).$ (3.28)
@@
$\left\|\bm{\Phi}^{m+1,l}-\bm{\Phi}^{m+1,l}\right\|_{L^{2}}\leqslant\eta_{1}.$ (3.30)
where $\eta_{1}$ is the small error tolerance, and the norm is $L^{2}$ norm. Let $\bm{\Phi}^{m+1,\ast}$ be the convergent solution of the Newton method, then we update $\bm{\Phi}^{m+1}=\bm{\Phi}^{m+1,\ast}$. Finally, the electric field $-\nabla\bm{\Phi}^{m+1}$
can be approximated using the first-order central difference scheme, i.e.,
$-\nabla\phi_{j}\approx-\frac{\phi_{j+1}-\phi_{j-1}}{2h},\qquad j=1,2,\dots,N_{h}.$ (3.31)
@@
Velocities, positions of particles $\{\left(X_{k}^{m},V_{k}^{m}\right)\}_{k=1}^{N_{\text{total}}}$ and macroscopic quantities $\bm{n}^{m}$, $\bm{J}^{m}$, $\bm{S}^{m}$ Compute $\bm{\Phi}^{m+1}$ by the reformulated Poisson equation using (3.23) with Newton’s method Compute electric field $-\nabla\bm{\Phi}^{m+1}$ through (3.31) Update velocities and positions of particles $\{\left(X_{k}^{m+1},V_{k}^{m+1}\right)\}_{k=1}^{N_{\text{total}}}$ by the ODE system (3.32) Compute density $\bm{n}^{m+1}$, current density $\bm{J}^{m+1}$ and momentum $\bm{S}^{m+1}$ by (3.4a)–(3.4c)
Figure 1: Flowchart of our AP-PIC method for solving the VPME system (3.15)–(3.16).
Remark 3.1.
Here, we provide an alternative approach to handle the nonlinear term $e^{\phi^{m+1}}$ in (3.22), which does not require the Newton’s iteration method. The idea of constructing a penalty term is inspired from [18] which studies the Boltzmann equation with stiffness. Consider
@@
$(A^{m}+I_{N_{h}\times N_{h}})\bm{\Phi}^{m+1}=\bm{b}^{m}+(\bm{\Phi}^{m}-e^{\bm{\Phi^{m}}}).$ (3.36)
@@
where $I$ is the unit matrix and $\bm{b}^{m}$ is defined in (3.24c). As $\varepsilon\to 0$, the scheme (3.36) automatically becomes a consistent discretization of the quasineutral limit equation (3.18), if one applying the same approximation for $e^{\phi^{m+1}}$ in the limit system. Therefore, this new explicit scheme also satisfies the asymptotic-preserving property. In the first numerical test, we compute the reformulated Poisson by this explicit scheme (3.36) and compare it with the implicit method using the Newton’s iteration. This explicit scheme has shown advantages since the iteration is avoided, and is computationally stable in time. To further explore this explicit scheme for the reformulated Poisson in the PIC framework and build higher-order AP method remains a future task.
4 A bi-fidelity method for model with random parameters

### v2

# A bi-fidelity method for the uncertain Vlasov-Poisson system near quasineutrality in an asymptotic-preserving particle-in-cell framework

@@
Abstract.
In this paper, we study the Vlasov-Poisson system with massless electrons (VPME) near quasineutrality and with uncertainties. Based on the idea of reformulation on the Poisson equation by [P. Degond et.al., Journal of Computational Physics, 229 (16), 2010, pp. 5630–5652], we first consider the deterministic problem and develop an efficient asymptotic preserving particle-in-cell (AP-PIC) method to capture the quasineutral limit numerically, without resolving the discretizations subject to the small Debye length in plasma. The main challenge and difference compared to previous related works is that we consider the nonlinear Poisson in the VPME system, which contains $e^{\phi}$ (with $\phi$ being the electric potential) and provide an explicit scheme. In the second part, we extend to study the uncertainty quantification (UQ) problem and develop an efficient bi-fidelity method for solving the VPME system with multidimensional random parameters, by choosing the Euler-Poisson equation as the low-fidelity model. Several numerical experiments are shown to demonstrate the asymptotic-preserving property of our deterministic solver and the effectiveness of our bi-fidelity method for solving the model with random uncertainties.
Key words and phrases:
Vlasov-Poisson near quasineutrality, uncertainty quantification, particle-in-cell, asymptotic-preserving, stochastic collocation, bi-fidelity method
1991 Mathematics Subject Classification:
65Y20, 35J05, 82C40, 82D10
Submitted to the editors January 16, 2025.
1. Introduction
@@
$\text{(VPME)}:\left\{\begin{array}[]{l}\displaystyle\partial_{t}f+\mathbf{v}\cdot\nabla_{\mathbf{x}}f-\nabla_{\mathbf{x}}\phi\cdot\nabla_{\mathbf{v}}f=0,\\[6.0pt] -\varepsilon^{2}\Delta_{\mathbf{x}}\phi=\int f\mathrm{d}\mathbf{v}-e^{\phi},\\[6.0pt] f(0,\mathbf{x},\mathbf{v})=f_{0},\,\int f_{0}\,\mathrm{d}\mathbf{v}\mathrm{d}\mathbf{x}=1.\end{array}\right$. (1.1)
Here $\varepsilon$ is the dimensionless parameter characterized by the Debye length and $f(t,\mathbf{x},\mathbf{v})$ represents the distribution of ions. This VPME system has been used in plasma physics to model ion plasma, study the formation of ion-acoustic shocks [49] and the expansion of plasma into vacuum [44], among other applications. We refer to [29] for more detailed introduction.
In real applications of plasma physics, these equations often exhibit different scales, ranging from the microscopic dynamics of individual particles to the macroscopic behavior of the plasma as a whole. When the scaled Debye length $\varepsilon\ll 1$, the plasma is called quasineutral: since the scale of charge separation is very small, the plasma appears to be neutral at the scale of observation. Quasineutrality is a common and important phenomenon for plasmas. In [11] the author emphasizes quasineutrality as one of the key properties that distinguish plasmas from ionised gases, we also refer to [27] and other theoretical works [7, 8, 9, 15, 22, 23, 24, 25, 26, 30, 31] which give a review of recent developments on for the Vlasov-type equations near quasineutrality.
@@
Numerical computation for the VPME system (1.1), that is under study in this work, seems underdeveloped. For the linear Poisson equation, a key idea was first proposed by P. Degond et al. [12], which is to reformulate the system to a new one that remains uniformly elliptic thus can capture the quasineutral limit of the system. Later in [16], the authors developed an asymptotic-preserving PIC method for the Vlasov-Poisson system that can capture the quasineutral limit at the discrete level. Previous works on AP methods for the quasineutral limit have been devoted to the Euler-Poisson model [13, 14, 17] and the Vlasov-Poisson problem [3]. In this work, we will develop an efficient AP-PIC method for solving the VPME system (1.1). For the deterministic problem, the main difficulty compared to the previous work [16] is that we consider the nonlinear Poisson equation with $e^{\phi}$ on the right-hand-side, thus the derivation of the quasineutral limit and the design of our discretized scheme is novel and new. Moreover, we provide a stable, explicit scheme that allows us to solve the reformulated Poisson equation efficiently.
@@
In industrial applications of plasma, it is important to study the uncertainty quantification (UQ) problems since there are many sources of uncertainties in the model. In the second part of this paper, we will extend to study UQ problems for the VPME system and consider the non-intrusive stochastic collocation (SC) method [1, 5, 46, 54] in a bi-fidelity framework. One key challenge of SC approaches is the computational cost, as they require repetitive implementations of the deterministic solver. Fortunately, in practice we can usually find some approximated low-fidelity models that are computationally cheaper but can still characterize behaviour of the complex model in the random space. The multi-fidelity algorithm [45, 55] is to combine the computational efficiency of low-fidelity models with the high accuracy of high-fidelity models, to construct an accurate surrogate for the high-fidelity model at a significantly reduced simulation cost. The feasibility of multi-fidelity methods and empirical error estimations to predict whether the multi-fidelity approach works or not for practical problems, has been discussed in [21]. Regarding UQ for kinetic equations with random parameters, one can refer to a series of works using the bi-fidelity methods [4, 41, 42, 48, 40].
The novelty of this work is two-fold: (i) The deterministic solver for the VPME system, namely the AP-PIC method developed and numerically implemented is new, inspired by yet different from the earlier work [47] where the macroscopic equations are studied. Our designed numerical scheme in the PIC framework satisfies the asymptotic-preserving property and can capture the quasineutral limit on the discrete level. (ii) This is the first work on studying the VPME system with multi-dimensional random uncertainties, by employing the multi-fidelity approach. We point out that in each problem setting, the choice of the low-fidelity model can be very different and is not unique, which directly affects the accuracy and efficiency of the bi-fidelity approximation. In the second part of our work, the novel part of solving UQ problems is to choose the macroscopic Euler system as the low-fidelity solver, while using the AP-PIC method developed in (i) as a robust high-fidelity solver.
@@
The rest of the paper is organized as follows. In Section 2, we first give an introduction of the Vlasov-Poisson system with massless electron and its related macroscopic equations. In Section 3, we begin with reviewing the classical PIC method, then derive the reformulated Poisson equation that leads to the design of our fully discretized AP-PIC scheme, with a summary of the method shown in Section 1. In Section 4, we introduce the bi-fidelity method and how we choose the high-fidelity and low-fidelity model (Euler-Poisson) in solving the VPME system with uncertainties. Several numerical examples, including the deterministic and problem with random parameters have been shown in Section 5, which demonstrate the effectiveness and AP property of our solver for the VPME system, in addition to accuracy and efficiency of our designed bi-fidelity method for the corresponding UQ problems. Finally, we conclude the paper and mention some future work.
2. Models
@@
(VPME): $\partial_{t}f+\mathbf{v}\cdot\nabla_{\mathbf{x}}f-\nabla_{\mathbf{x}}\phi\cdot\nabla_{\mathbf{v}}f=0,$ (2.1a)
(VPME): $-\varepsilon^{2}\Delta_{\mathbf{x}}\phi=\int f\mathrm{d}\mathbf{v}-e^{\phi},$ (2.1b)
(VPME): $f(0,\mathbf{x},\mathbf{v})=f_{0},\,\int f_{0}\,\mathrm{d}\mathbf{v}\mathrm{d}\mathbf{x}=1.$ (2.1c)
where $f(t,\mathbf{x},\mathbf{v})$ is interpreted as the distribution of ions, at position $\mathbf{x}$ moving with velocity $\mathbf{v}$. Here $\phi$ is the electric potential that is governed by the Poisson equation, and $\varepsilon$ is the dimensionless parameter characterized by the Debye length. More details of derivation of the model can be found in [30].
@@
$f(t,\mathbf{x},\mathbf{v})=n(t,\mathbf{x})\delta(\mathbf{v}-\mathbf{u}(t,\mathbf{x})),$ (2.2)
with $\delta$ being the Dirac delta function. Here, $n(t,\mathbf{x})$ is the ion density and $u(t,\mathbf{x})$ is the ion velocity given by
$n(t,\mathbf{x})=\int f(t,\mathbf{x},\mathbf{v})\mathrm{d}\mathbf{v},\quad\mathbf{u}(t,\mathbf{x})=\frac{1}{n(t,\mathbf{x})}\int\mathbf{v}f(t,\mathbf{x},\mathbf{v})\mathrm{d}\mathbf{v}.$ (2.3)
Then one derives isothermal compressible Euler system (ICE) by taking the first and second moments in velocity:
$\text{(ICE): }\left\{\begin{array}[]{l}\partial_{t}n+\nabla_{\mathbf{x}}\cdot(n\mathbf{u})=0,\\[6.0pt] \partial_{t}(n\mathbf{u})+\nabla_{\mathbf{x}}\cdot(n\mathbf{u}\otimes\mathbf{u})=-\nabla_{\mathbf{x}}n.\end{array}\right$. (2.4)
For the VPME system (2.1c), periodic boundary conditions are assumed for the distribution $f(t,\mathbf{x},\mathbf{v})$ and potential $\phi$. In the one-dimensional case, the boundary conditions are given by
@@
3. AP-PIC method for VPME system
3.1. Classical PIC method
@@
3.2. Derivation of the reformulated Poisson equation
@@
$\nabla\cdot\left((\varepsilon^{2}\partial_{tt}+n_{i})\nabla\phi\right)+\partial_{tt}n_{e}=\nabla^{2}:S_{i},$ (3.11)
@@
$\begin{array}[]{l}-\varepsilon^{2}\Delta\phi_{0}=\left(n_{i}-n_{e}\right)_{0},\\[4.0pt] -\varepsilon^{2}\Delta\phi_{0}^{\prime}=\left(n_{i}-n_{e}\right)_{0}^{\prime}=-\nabla\cdot\left((nu)_{i}-(nu)_{e}\right)_{0},\end{array}$ (3.12)
@@
$\nabla\cdot\left((\varepsilon^{2}\partial_{tt}+n_{i})\nabla\phi\right)+\partial_{tt}e^{\phi}=\nabla^{2}:S_{i}.$ (3.14)
@@
where $\phi$ need to satisfy the initial data (3.12). As $\varepsilon\to 0$, the quasineutral Vlasov system becomes
@@
We note that although the original Poisson equation (3.13) and the reformulated Poisson equation (3.14) are equivalent, they are not equally suited in the quasi-neutral limit. In fact, the quasi-neutral potential equation (3.18) automatically serves as the formal limit $\varepsilon\to 0$ of the reformulated Poisson equation (3.14), but not of the original Poisson equation (3.13) which gives $n_{i}=e^{\phi}$ as $\varepsilon\to 0$. Therefore, in order to design asymptotic-preserving schemes, it is necessary to base on the reformulated system (3.17)–(3.18) for numerical discretization. We will introduce our AP scheme in the PIC framework in the next section.
3.3. The AP-PIC method
@@
$\nabla\cdot\Big((\Delta t)^{2}n_{i}^{m}\nabla\phi^{m+1}+\varepsilon^{2}\nabla\phi^{m+1}\Big)+e^{\phi^{m+1}}\\ =(\Delta t)^{2}\nabla^{2}:S_{i}^{m}+2n_{i}^{m}-n_{i}^{m-1}.$ (3.20)
@@
$\nabla\cdot\Big((\Delta t)^{2}n_{i}^{m}\nabla\phi^{m+1}\Big)+e^{\phi^{m+1}}=(\Delta t)^{2}\nabla^{2}:S_{i}^{m}+2n_{i}^{m}-n_{i}^{m-1}.$ (3.21)
@@
Consider the right-hand side of (3.20), the term $n_{i}^{m}-n_{i}^{m-1}$ can be replaced by $\nabla(nu)_{i}^{m}$, we derive an equivalent form shown as
$\nabla\cdot\Big((\Delta t)^{2}n_{i}^{m}\nabla\phi^{m+1}+\varepsilon^{2}\nabla\phi^{m+1}\Big)+e^{\phi^{m+1}}\\ =(\Delta t)^{2}\nabla^{2}:S_{i}^{m}+n_{i}^{m}-\nabla(nu)_{i}^{m}.$ (3.22)
@@
$\begin{array}[]{l}\left(D^{+}g\right)_{j}=\frac{g_{j+1}-g_{j}}{h},\\[4.0pt] \left(D^{-}g\right)_{j}=\frac{g_{j}-g_{j-1}}{h},\\[4.0pt] \left(\Delta g\right)_{j}=\left(D^{+}D^{-}g\right)_{j}=\left(D^{-}D^{+}g\right)_{j}=\frac{g_{j+1}-2g_{j}+g_{j-1}}{h^{2}}.\end{array}$
@@
$F(\mathbf{\Phi}^{m+1})=A^{m}\mathbf{\Phi}^{m+1}+\mathrm{diag}(e^{\mathbf{\Phi}^{m+1}})-\mathbf{b}^{m}=0,$ (3.24a)
$\mathbf{\Phi}^{m+1}=\left[\phi^{m+1}_{1},\phi^{m+1}_{2},\dots,\phi^{m+1}_{N_{h}}\right]^{\top},$ (3.24b)
@@
The Newton iteration needs the Jacobian of $F(\mathbf{\Phi}^{m+1})$ defined in (3.24a), which is
$J(\mathbf{\Phi}^{m+1})=A^{m}+\mathrm{diag}(e^{\mathbf{\Phi}^{m+1}}).$ (3.28)
@@
$\left\|\mathbf{\Phi}^{m+1,l}-\mathbf{\Phi}^{m+1,l}\right\|_{L^{2}}\leq\eta_{1}.$ (3.30)
where $\eta_{1}$ is the small error tolerance, and the norm is $L^{2}$ norm. Let $\mathbf{\Phi}^{m+1,\ast}$ be the convergent solution of the Newton method, then we update $\mathbf{\Phi}^{m+1}=\mathbf{\Phi}^{m+1,\ast}$. Finally, the electric field $\nabla\mathbf{\Phi}^{m+1}$ can be approximated using the first-order central difference scheme, i.e.,
$\nabla\phi_{j}\approx-\frac{\phi_{j+1}-\phi_{j-1}}{2h},\qquad j=1,2,\dots,N_{h}.$ (3.31)
@@
Velocities, positions of particles $\{\left(X_{k}^{m},V_{k}^{m}\right)\}_{k=1}^{N_{\text{total}}}$ and macroscopic quantities $\mathbf{n}^{m}$, $\mathbf{J}^{m}$, $\mathbf{S}^{m}$ Compute $\mathbf{\Phi}^{m+1}$ by the reformulated Poisson equation using (3.23) with Newton’s method Compute electric field $\nabla\mathbf{\Phi}^{m+1}$ through (3.31) Update velocities and positions of particles $\{\left(X_{k}^{m+1},V_{k}^{m+1}\right)\}_{k=1}^{N_{\text{total}}}$ by the ODE system (3.32) Compute density $\mathbf{n}^{m+1}$, current density $\mathbf{J}^{m+1}$ and momentum $\mathbf{S}^{m+1}$ by (3.4a)–(3.4c)
Figure 1. Flowchart of our AP-PIC method for solving the VPME system (3.15)–(3.16).
{rmrk}
Here, we provide an alternative approach to handle the nonlinear term $e^{\phi^{m+1}}$ in (3.22), which does not require the Newton’s iteration method. The idea of constructing a penalty term is inspired from [37]. Consider
@@
$(A^{m}+I_{N_{h}\times N_{h}})\mathbf{\Phi}^{m+1}=\mathbf{b}^{m}+(\mathbf{\Phi}^{m}-e^{\mathbf{\Phi^{m}}}).$ (3.36)
@@
where $I$ is the unit matrix and $\mathbf{b}^{m}$ is defined in (3.24c). As $\varepsilon\to 0$, the scheme (3.36) automatically becomes a consistent discretization of the quasineutral limit equation (3.18), if one applying the same approximation for $e^{\phi^{m+1}}$ in the limit system. Therefore, this new explicit scheme also satisfies the asymptotic-preserving property. In the first numerical test, we compute the reformulated Poisson by this explicit scheme (3.36) and compare it with the implicit method using the Newton’s iteration. This explicit scheme has shown advantages since the iteration is avoided, yet satisfies the AP property. To further explore this explicit scheme for the reformulated Poisson in the PIC framework and build higher-order AP method remains a future task.
4. A bi-fidelity method for model with random parameters